Director's Time

Observers: B.Rothberg, D. Thompson, B.Frye (Tucson)

ISA: B.Rothberg (Tucson)
Telescope Operator: D.G.Huerta (LBTO)
AO Support: JCGuerra
Telescope Support: M.Wagner (LBTO)
Software Support: S.Hooper
Instrument Support: J.Morris (LBTO)
Engineering Support: DAshby

Summary

Details

UT 12:11 Preparing to open

UT 12:37 Starting skyflats - Moving to Blank field CAblank7

UT 12:41 -LBC killed and restarted itself on its own after sending skyflat test OB 2nd time

From log:

2014/12/24 00:40:49.317086 S R TRACKERS powerdump: 192.168.59.114 outlets owners 00000000 last access 1 sec ago

2014/12/24 00:40:49.317149 N - accessing /lbccontrol/current/conf/redchannel.conf configuration file completed

2014/12/24 00:40:49.317327 N - waiting for commands ...

2014/12/24 00:40:49.317619 N - UI update with message >< at 1419381649.317354

2014/12/24 00:40:49.317893 S - read INTEGER parameter >TelescopeStatus< with value 1

2014/12/24 00:40:49.317926 S - read INTEGER parameter >BLUEChannelStatus< with value 31

2014/12/24 00:40:49.317944 S - read INTEGER parameter >REDChannelStatus< with value 31

2014/12/24 00:40:49.317962 S - read INTEGER parameter >LEFTConnectedStatus< with value 1

2014/12/24 00:40:49.317981 S - read INTEGER parameter >RIGHTConnectedStatus< with value 1

2014/12/24 00:40:49.318006 N - TCSQueryParameters initializing TCS

2014/12/24 00:40:49.318152 N - CrashRecovery initiating TCS reconnect

2014/12/24 00:40:49.318353 N - CrashRecovery initiating HK recovery

2014/12/24 00:40:49.324128 N - UI update with message >LBC has restarted after a crash/kill and will automatically try to: - restore telescope connection if already open - restore the previous systems power status< at 1419381649.318113

UT 12:44 - Running skyflat test again for Uspec and r-sloan - flats still saturated

2014/12/24 00:46:34.436806 N - CrashRecovery initiating HK recovery

2014/12/24 00:46:34.442757 N - UI update with message >LBC has restarted after a crash/kill and will automatically try to: - restore telescope connection if already open - restore the previous systems power status< at 1419381994.436576

UT 12:47 - LBC killed/restarted itself when OB "play" button was pressed

UT 12:48 - LBC is restarting - rerunning Skyflatest for Uspec and r-sloan - if have to run again to check counts will reload and see if lbc crashes again

UT 12:50 - Counts OK for Uspec and r-sloan - loaded flat OB and lbc crashed again

UT 12:52 - lbc restart completed - called Kellee about the crashes

UT 12:54 - reset skyflat OB for Uspec and r-sloan

UT 12:55 - LBC crashed again, lbckill/lbcstart issued automagically after 1 exposure taken on each camera - counts 7k Uspec, 13k r-sloan

UT 01:00 - Kellee checked and backed out the new TCS part of 2.1 - TCS is back to the C-interface - reran 5pt dither, counts are low 3-4k for each

UT 01:05 - Hit the stop to stop the 4x scale - but not sure the GUI actually stopped the current set - they seemed to complete

UT 01:08 - scaled by 5x but received pop-up error "OB upload error: 4"

UT 01:13 - running sloan g, I-bessell skyflat test - scaled by 4x from script

UT 01:15 - running scaled by 10x from script -test

UT 01:17 - running test scaled by 14x

UT 01:20 - running full sky flat script for g-sloan and I-bessell scaled by 20x - getting 3k counts in g-sloan and 10k counts in I

UT 01:39 - attempting dofpia - failed - no good pupils found - new lookup tables didn't load after hub movement - right side should be -0.3 Z but still says -0.8

UT 01:54 - cleared Active Optics, reauthorized, went back to target - retrying dofpia

UT 02:03 - dofpia seemed unable to find file in /newdata, paused, then it found itUT

======================================================

UT 02:15 - Running Gallddi 1 for Uspec, Ubessell, and Ibessell at airmass 1.011 - copointing OK and focus OK (though took some time)

UT 02:21 - Std field done - moving to OB1 to do DOFPIA

======================================================

UT 02:21 - D.Thompson completed Darks for B.Frye DDT program #0001-0002 4s; #0003-0004 12s; #0005-0010 10x4s; #0011-0020 3x30s

======================================================

UT 02:29 - Executing first loop of CV OB1-OB7

UT 02:30 - Guider image for red looks elongated, blue looks OK - OB1

UT 02:47 - seeing check on OB2 field - 2.1" Blue, 2.0" Red

UT 03:11 - Loop 1 done, moving to OB1 to run DOFPIA

UT 03:13 - Running DOFPIA at OB1 field

UT 03:18 - Starting Loop 2 - Red 1.67" Blue 1.79" seeing estimate from DOFPIA

UT 03:55 - LBC crashed/restarted in the middle of OB7 - crash was during exposure - exposures lost for OB7 - from python script: LBC: RUNNING 35/240 g-SLOAN exposure 40/240 r-SLOAN exposure 1 1

UT 03:57 - Manually reloaded OB7 - will go to OB1 again when complete and run DOFPIA - blue side guiding looked faint DIMM 1.5-2.0"

UT 04:05 - DOFPIA

UT 04:07 - DOFPIA didn't find file in /newdata - paused and waited for confirmation file was present

UT 04:10 - 040443 red shows double images, 040439 blue shows double images

UT 04:12 - The stop button didn't stop the OB1 that was loaded - 040439 and 040443 were taken while the DOFPIA was running - stopped the OB again and restarted DOFPIA - not sure if the STOP didn't work or I didn't click hard enough - Kellee will check logs.

UT 04:24 - DOFPIA completed - Kellee is reverting to 2.01 LBC software

UT 04:28 - Starting Loop 3

UT 05:12 - Finished CV program

======================================================

UT 05:12 - Moving to Galaddi 1 Std field for DOFPIA

UT 05:13 - Running DOFPIA

UT 05:16 - DIMM 1.7" - DOFPIA reports 1.4" red and 1.1" blue

UT 05:18 - Galaddi-1 Std Field Uspec,U,I-Bessell - airmass 1.45

UT 05:27 - Moving to NGC 890 - Science Field BR

UT 05:29 - Running DOFPIA since moved from 1.45 to 1.09 airmass - seeing 0.88" red 1.02" blue from DOFPIA

UT 05:34 - Starting NGC 890 OB

UT 05:45 - Cancelled OB - after 3 exposures (1 Uspec, 2 I) - Drives Dropped

UT 06:11 - Elevation and Azimuth drives not responding - attempting to contact people to help

UT 06:53 - Drive problem solved

UT 06:55 - Slewing back to NGC 890 - will do DOFPIA, clear active optics

UT 06:59 - Running DOFPIA

UT 07:07 - Restarting NGC 890 OB

UT 07:29 - DIMM 1.2

UT 08:06 - stopped at 23/30 - at airmass 1.8

======================================================

UT 08:07 - Reconfiguring for LUCI on SX

UT 08:31 - Reconfigured for LUCI

UT 08:41 - Collimating in SN field - B. Frye target (official handover is at UT 08:51)

UT 08:44 - GCS 1.1" at airmass 1.72

UT 09:00 - Still attempting to setup the AO - D. Miller working on changing to a brighter guidestar - resending AO presets

UT 09:16 - LUCI panels showing temperature warning on shield - J. Morris checking

UT 09:26 - filter wheel was blocking acquisition camera - awaiting word on LUCI shield temp

UT 09:33 - reference star appears to be 16th mag not 14th mag - affected setting up AO

UT 10:05 - started in tip tilt mode - script started

UT 10:09 - loop opened - script stopped at 2nd exposure

UT 10:25 - loop closed - script 1 started

UT 10:27 - offset failed

UT 10:37 - restarted script again

UT 10:43 - created script with smaller offsets - resetting to 0,0 position

UT 10:47 - LUCI Shield temp warning no longer in yellow - back to 61K

UT 10:53 - short offset script started (loops closed)

UT 10:55 - offset worked

UT 11:41 - stopped after 27min to attempt full AO corrections - just got out of seeing bubble - GCS 0.7"

UT 11:42 - script 2b started

UT 11:45 - #0045 loop opened - no corrections

UT 11:47 - restarting again to do full AO again

UT 11:48 - failed to close loop - too unstable for full AO

UT 12:27 - loop opened - 15/30 completed

UT 12:36 - script started - loop closed

UT 12:54 - loop opened - stopping script - completed 9/30

UT 13:01 - closed loop - running script

UT 13:07 - loop opened - completed 2/30

UT 13:18 - closed loop - script started

UT 13:35 - loop opened - 8/30 completed

UT 13:50 - Taking Ks Twilight Flats

UT 14:08 - Finished B. Frye Program
Topic revision: r3 - 24 Dec 2014, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback